# RGS-SLAM: Robust Gaussian Splatting SLAM with One-Shot Dense Initialization

### v1

@@ 1 Introduction
• Single-Step Dense Initialization. A one-shot triangulation replaces residual-driven densification within the standard GS-SLAM pipeline MonoGS [matsuki2024monogs], enabling stationary optimization and 20% faster convergence.
@@ 2 Related Work
3D Gaussian Splatting and Densification. 3D Gaussian splatting (3DGS) enables real-time view synthesis with anisotropic splats and a visibility-aware renderer, yet most systems rely on residual-driven densification [kerbl2023gaussian]. We instead seed a fixed topology from dense multi-view correspondences and then refine only splat parameters.
Gaussian Splats for SLAM. SLAM systems mapping with Gaussians include GS-SLAM, MonoGS, SplaTAM, and Gauss-SLAM [yan2024gsslam, matsuki2024monogs, keetha2024splatam]. They rely on densification, causing early non-stationarity. Our training-free dense seed removes this stage and drops into MonoGS with minimal modifications.
Differentiable Rendering and Real-Time SLAM. Photo-SLAM, GLORIE-SLAM, and Point-SLAM couple differentiable rendering with pose optimization for fast updates via analytic/lightweight gradients [huang2024photoslam, zhang2024glorieslam, sandstrom2023pointslam]. We retain this in a Gaussian renderer and use a stationary initialization so early steps do not change topology.
Feature Matching and Dense Correspondence. SuperPoint/SuperGlue remain strong baselines [detone2018superpoint, sarlin2020superglue]. Detector-free transformers (LoFTR, LightGlue) extend coverage in low-texture regions, and dense matchers (DKM) provide broad two-view coverage with confidence for geometry [sun2021loftr, lindenberger2023lightglue, edstedt2023dkm]. We aggregate confidence-weighted dense correspondences over a short keyframe window into an explicit Gaussian seed.
Initialization and Training-Free Priors. SfM and multi-view stereo stabilize early optimization via geometric priors and learned depth [schonberger2016sfm, yao2018mvsnet]. In Gaussian splatting, schedules and regularizers typically retain densification. Our correspondence-to-Gaussian initialization follows training-free priors and yields a stationary objective from the start.
@@ 3.4 Dense Feature Matching
We extract dense visual descriptors using DINOv3 [Simeoni2025dinov3], which provide semantically consistent features across views. These descriptors are used to establish multi-view dense correspondences, replacing the residual-driven densification process in GS-SLAM. A confidence-aware inlier classifier is then applied to filter unreliable matches, ensuring stable multi-view geometry. Finally, a one-shot triangulation is performed to initialize a uniformly distributed set of 3D Gaussian seeds.
@@ 4.1 Experimental Setup
Baseline Methods. We compare with iMAP [sucar2021imap], NICE-SLAM [zhu2022niceslam], Vox-Fusion [yang2022voxfusion], ESLAM [johari2023eslam], Point-SLAM [sandstrom2023pointslam], Co-SLAM [wang2023coslam], SplaTAM [keetha2024splatam], Gauss-SLAM [yan2024gsslam], and MonoGS [matsuki2024monogs]. We also include SNI-SLAM [zhu2024snislam] for reconstruction and Photo-SLAM [huang2024photoslam], GLORIE-SLAM [zhang2024glorieslam], RK-SLAM [ma2025rkslam] for rendering. RGB-D–only methods run in RGB-D, with monocular results reported only when supported. Hyperparameters follow official documented defaults on the same splits.
@@ 4.3 Localization Accuracy
iMAP [sucar2021imap] 3.12 2.54 2.31 1.69 1.03 3.99 4.05 1.93 2.58
NICE-SLAM [zhu2022niceslam] 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.07
Vox-Fusion [yang2022voxfusion] 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94 3.09
ESLAM [johari2023eslam] 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.63
Point-SLAM [sandstrom2023pointslam] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.53
MonoGS [matsuki2024monogs] 0.62 0.62 0.77 0.44 0.52 0.23 0.62 2.25 0.76
@@
iMAP [sucar2021imap] 4.90 2.00 5.80 4.23
NICE-SLAM [zhu2022niceslam] 4.26 6.19 3.87 4.77
DI-Fusion [huang2021difusion] 4.40 2.00 5.80 4.07
Vox-Fusion [yang2022voxfusion] 3.52 1.49 26.01 10.34
ESLAM [johari2023eslam] 2.47 1.11 2.42 2.00
Co-SLAM [wang2023coslam] 2.40 1.70 2.40 2.17
Point-SLAM [sandstrom2023pointslam] 4.34 1.31 3.48 3.04
MonoGS [matsuki2024monogs] 1.50 1.44 1.49 1.47
@@ 4.4 Rendering Quality and Throughput
We report fidelity and throughput in Table 4 and 5. On Replica, our initializer averages 925 FPS, exceeding MonoGS (769 FPS) while maintaining competitive PSNR [hore2010psnr], SSIM [wang2004ssim], and LPIPS [zhang2018lpips] across room0–2 and office0–4. On TUM, the system runs in real time (2.5–3.2 FPS) with PSNR/SSIM comparable to SplaTAM, PhotoSLAM, and GLORIE-SLAM, and low LPIPS. The throughput gain arises from the keyframe-triggered single-step den-
@@ 4.5 Reconstruction Fidelity
iMAP [sucar2021imap] 3.624 4.934 80.515
NICE-SLAM [zhu2022niceslam] 2.373 2.645 91.137
Vox-Fusion [yang2022voxfusion] 1.882 2.563 90.936
Co-SLAM [wang2023coslam] 2.104 2.082 93.435
ESLAM [johari2023eslam] 2.082 1.754 96.427
SNI-SLAM [zhu2024snislam] 1.942 1.702 96.624

### v2

@@ 1 Introduction
• Single-Step Dense Initialization. A one-shot triangulation replaces residual-driven densification within the standard GS-SLAM pipeline MonoGS [11], enabling stationary optimization and 20% faster convergence.
@@ 2 Related Work
3D Gaussian Splatting and Densification. 3D Gaussian splatting (3DGS) enables real-time view synthesis with anisotropic splats and a visibility-aware renderer, yet most systems rely on residual-driven densification [8]. We instead seed a fixed topology from dense multi-view correspondences and then refine only splat parameters.
Gaussian Splats for SLAM. SLAM systems mapping with Gaussians include GS-SLAM, MonoGS, SplaTAM, and Gauss-SLAM [20, 11, 7]. They rely on densification, causing early non-stationarity. Our training-free dense seed removes this stage and drops into MonoGS with minimal modifications.
Differentiable Rendering and Real-Time SLAM. Photo-SLAM, GLORIE-SLAM, and Point-SLAM couple differentiable rendering with pose optimization for fast updates via analytic/lightweight gradients [4, 23, 12]. We retain this in a Gaussian renderer and use a stationary initialization so early steps do not change topology.
Feature Matching and Dense Correspondence. SuperPoint/SuperGlue remain strong baselines [1, 13]. Detector-free transformers (LoFTR, LightGlue) extend coverage in low-texture regions, and dense matchers (DKM) provide broad two-view coverage with confidence for geometry [17, 9, 2]. We aggregate confidence-weighted dense correspondences over a short keyframe window into an explicit Gaussian seed.
Initialization and Training-Free Priors. SfM and multi-view stereo stabilize early optimization via geometric priors and learned depth [14, 22]. In Gaussian splatting, schedules and regularizers typically retain densification. Our correspondence-to-Gaussian initialization follows training-free priors and yields a stationary objective from the start.
@@ 3.4 Dense Feature Matching
We extract dense visual descriptors using DINOv3 [15], which provide semantically consistent features across views. These descriptors are used to establish multi-view dense correspondences, replacing the residual-driven densification process in GS-SLAM. A confidence-aware inlier classifier is then applied to filter unreliable matches, ensuring stable multi-view geometry. Finally, a one-shot triangulation is performed to initialize a uniformly distributed set of 3D Gaussian seeds.
@@ 4.1 Experimental Setup
Baseline Methods. We compare with iMAP [16], NICE-SLAM [26], Vox-Fusion [21], ESLAM [6], Point-SLAM [12], Co-SLAM [18], SplaTAM [7], Gauss-SLAM [20], and MonoGS [11]. We also include SNI-SLAM [25] for reconstruction and Photo-SLAM [4], GLORIE-SLAM [23], RK-SLAM [10] for rendering. RGB-D–only methods run in RGB-D, with monocular results reported only when supported. Hyperparameters follow official documented defaults on the same splits.
@@ 4.3 Localization Accuracy
iMAP [16] 3.12 2.54 2.31 1.69 1.03 3.99 4.05 1.93 2.58
NICE-SLAM [26] 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.07
Vox-Fusion [21] 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94 3.09
ESLAM [6] 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.63
Point-SLAM [12] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.53
MonoGS [11] 0.62 0.62 0.77 0.44 0.52 0.23 0.62 2.25 0.76
@@
iMAP [16] 4.90 2.00 5.80 4.23
NICE-SLAM [26] 4.26 6.19 3.87 4.77
DI-Fusion [5] 4.40 2.00 5.80 4.07
Vox-Fusion [21] 3.52 1.49 26.01 10.34
ESLAM [6] 2.47 1.11 2.42 2.00
Co-SLAM [18] 2.40 1.70 2.40 2.17
Point-SLAM [12] 4.34 1.31 3.48 3.04
MonoGS [11] 1.50 1.44 1.49 1.47
@@ 4.4 Rendering Quality and Throughput
We report fidelity and throughput in Table 4 and 5. On Replica, our initializer averages 925 FPS, exceeding MonoGS (769 FPS) while maintaining competitive PSNR [3], SSIM [19], and LPIPS [24] across room0–2 and office0–4. On TUM, the system runs in real time (2.5–3.2 FPS) with PSNR/SSIM comparable to SplaTAM, PhotoSLAM, and GLORIE-SLAM, and low LPIPS. The throughput gain arises from the keyframe-triggered single-step den-
@@ 4.5 Reconstruction Fidelity
iMAP [16] 3.624 4.934 80.515
NICE-SLAM [26] 2.373 2.645 91.137
Vox-Fusion [21] 1.882 2.563 90.936
Co-SLAM [18] 2.104 2.082 93.435
ESLAM [6] 2.082 1.754 96.427
SNI-SLAM [25] 1.942 1.702 96.624